# Characterizing NGC 6383: A study of pre-main sequence stars, mass segregation, and age using Gaia DR3 and 2MASS

### v1

@@ 2.1.1 Membership determination
To identify potential members of NGC 6383, we employed the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (Campello et al., 2013), known for its effectiveness in identifying clusters of varying densities and sizes without requiring a predetermined number of clusters (Hunt & Reffert, 2021). This density-based clustering method, which constructs a hierarchical representation of the data (McInnes et al., 2017), is particularly adept at handling different cluster shapes and sizes while distinguishing noise points. Its capacity to automatically determine the optimal cluster count makes HDBSCAN an ideal choice for this analysis, offering significant benefits over traditional clustering algorithms, especially in terms of robustness against noisy data and outliers.
Figure 4: Left panel: Histograms display parallax measurements for NGC 6383. Used Parallax in red represents measurements with fractional parallax errors less than 0.1, Parallax in orange includes all data, and Non-Used Parallax in green for sources with a fractional parallax error greater than 0.1. The blue curve shows a Gaussian fit to the used parallax data, with the mean parallax value $\mu_{\varpi}=0.908\pm 0.004\,\mathrm{mas}$ marked by the dashed green line. Middle panel: G magnitude versus parallax values, with data utilized for distance estimation in blue and discarded due to having a fractional parallax error greater than 0.1 in red. Grey error bars illustrate the uncertainty in parallax measurements, especially pronounced at fainter magnitudes. Right panel: Histograms of distances derived from the inverse of parallax to differentiate between used (red) and all measured (orange) parallaxes, including non-used (green). The central dashed blue line marks the mean sampled distance ($\mu_{d}=1.107\pm 0.035\,\mathrm{kpc}$), with its standard deviation area shaded in blue.
@@
After narrowing the data to a final subset of $15480$ sources, HDBSCAN was applied, focusing on proper motions as the primary clustering parameters.
@@
To refine our selection of hyperparameters for optimal cluster identification, we utilized a systematic approach. This involved iterative testing of various minimum cluster sizes to find the setup that enhances cluster size while ensuring high cluster separation strength, as indicated by the maximum $\lambda$ value. Our strategy aimed at selecting leaf clusters to generate numerous small, homogeneous groups, aiming to include the maximum number of stars without compromising the separation strength. While this approach risks increasing false positives, it is essential for capturing all potential cluster members, particularly for studying tidal sources. To mitigate this risk, we derived a pseudo-probability metric. This metric was calculated by determining the frequency with which a source was identified as a cluster member across multiple iterations, divided by the total number of iterations, and then multiplying this value by the HDBSCAN membership probability. Sources with a resulting probability greater than 0.5 were retained. The variation in cluster sizes and the threshold for minimum cluster size are depicted in Fig. 16. Additionally, we used a Euclidean metric for clustering based on proper motions, finding that while the Haversine metric was tested, it produced non-physical clusters.
@@
Finally, we used the AstroPy Sigma Clipping (Astropy Collaboration et al., 2013, 2018, 2022) utility to perform a $2\sigma$ clipping method around the mode in order to obtain the final likely members of the cluster.
@@ 2.1.2 Parallax and Distance Estimation
Distance estimation from parallax is challenging due to measurement errors, and transforming parallax to distance is not straightforward and involves significant uncertainties, deviating from the simple inverse of the parallax approach (Bailer-Jones, 2015; Bailer-Jones et al., 2021). A crucial aspect is choosing an appropriate prior distribution. When analyzing the parallax measurements of stars within an open cluster and assuming a simplified one-dimensional perspective, observing the distribution of parallax sources, we can assume that it follows a normal distribution. For accuracy, our analysis only includes members with fractional parallax errors under $0.1$.
Our hierarchical model begins by calculating an initial average distance $(\mu_{\mathrm{prior}})$ from the observed parallaxes. The cluster’s mean distance follows $\mathcal{U}(0.5\mu_{\mathrm{prior}},1.5\mu_{\mathrm{prior}})$ as a prior, with an assumed normal likelihood.
Figure 5: Radial density profile of NGC 6383 shown in logarithmic scale. The plot overlays the observed stellar density (blue points) against the King model (solid fuchsia line). Dashed red and dotted red lines mark the core radius ($R_{c}=1.944\,\text{arcmin}$) and the tidal radius ($R_{t}=40.713\,\text{arcmin}$), respectively. The dash-dot green line represents the $b$ parameter of the background level. The shaded area in blue indicates the $1\sigma$ uncertainty range of the King profile.
@@ 2.1.3 Proper motions
To estimate the distribution of cluster members’ proper motions, we model them using a two-dimensional normal distribution.
A normal prior distribution was assigned to the mean proper motions $\left[\overline{\mu_{\alpha}},\overline{\mu_{\delta}}\right]$, based on frequentist means and standard deviations. The standard deviations are modeled using a HalfNormal distribution, while the correlation coefficient between the proper motions is uniformly distributed over the interval $[-1,1]$.
We also estimated the mean projected velocity of the cluster, defined as the square root of the sum of squares of the proper motion components.
Figure 6: Spatial distribution of NGC 6383 sources overlaid on a DSS2-red image, centered on the calculated cluster center marked by a blue cross. The image details various structural and dynamical boundaries of the cluster as follows: the tidal radius $R_{t}$ is depicted with a dashed dark slate grey line, the core radius $R_{c}$ with a solid dark blue line, the half-light radius $R_{\mathrm{hl}}$ with a dash-dot magenta line, the half-mass radius $R_{\mathrm{hm}}$ with a solid yellow line, the Hill radius $R_{\mathrm{hill}}$ with a dashed orange line, and the gravitational bound radius $R_{\mathrm{bound}}$ with a solid red line. A scale bar indicating 5 parsecs is positioned at the bottom right to provide spatial reference relative to the cluster scale.
2.1.4 Center determination
To determine the cluster’s center, we used a weighted Kernel Density Estimation (KDE) from sci-kit learn (Pedregosa et al., 2011), assigning weights inversely proportional to the distance from the mean proper motion. Optimal KDE parameters were determined through Grid Search Cross-Validation within the package, exploring a range of bandwidths—from the mean positional error to the search cone radius—and all kernel types. The location of maximum density was designated as the cluster’s center.
2.1.5 Radial density profile
To determine structural parameters of the cluster, we utilized the King (1962) density profile.
The numerical density $\rho$, was computed by dividing the cluster area into concentric annuli, each containing an equal number of stars. The number of annuli, $K$, adheres to the equiprobable bin rule ($K=2n^{2/5}$), where $n$ is the star count within the cluster. Fitting King’s model to the density data allows for estimating the structural parameters’ posterior distributions.
Priors for the model parameters were set as follows: $b\sim\mathcal{U}(0,2\rho_{\mathrm{min}})$, $k\sim\mathcal{U}(0,2\rho_{\mathrm{max}})$, $R_{c}\sim\mathcal{U}(0,0.8R_{t})$, and $R_{t}\sim\mathcal{U}(R_{c},1.5T_{\mathrm{max}})$, where $T_{\mathrm{max}}$ represents the highest value among Hill radius and gravitational bound radius. This ensures the chosen parameter space for $R_{c}$ and $R_{t}$ is within astrophysical valid limits.
The Hill Radius $(R_{\mathrm{hill}})$ can be calculated following the methodology outlined by Carrera et al. (2019):
where $R_{\mathrm{gc}}$ denotes the galactocentric distance, $m_{c}$ the cluster mass from Hunt & Reffert (2024), with a value of $902.27~{}\mathrm{M_{\odot}}$, and $M_{\mathrm{gc}}$ the Galactic mass within the cluster’s orbit calculated as
$M_{\mathrm{gc}}=2\times 10^{8}\,M_{\odot}\left(\frac{R_{G}}{30\,\mathrm{pc}}\right)^{1.2}.$ (2)
$r_{\mathrm{bound}}=\left(\frac{GM_{\mathrm{total}}}{2(A-B)^{2}}\right)^{1/3},$ (3)
@@ 2.1.5 Radial density profile
Distance $(\varpi)$ $1.107\pm 0.035$ $\mathrm{kpc}$
Distance (D.M.) $1.211_{0.858}^{1.709}$ $\mathrm{kpc}$
Distance Modulus $10.416\pm 0.748$ $\mathrm{mag}$
Age $4\pm 1$ $\mathrm{Myr}$
Stellar formation range $1-6$ $\mathrm{Myr}$
Metallicity ($Z$) $0.023\pm 0.008$
@@
Absorption $(A_{V})$ $1.221\pm 0.241$ $\mathrm{mag}$
Galactocentric distance $R_{\mathrm{GC}}$ $7.197$ $\mathrm{kpc}$
Core Radius $(R_{c})$ $1.943\pm 0.1857$ $\mathrm{arcmin}$
Background $(b)$ $0.011\pm 0.0062$ $\mathrm{stars\,arcmin^{-2}}$
Tidal Radius $(R_{t})$ $40.712\pm 14.379$ $\mathrm{arcmin}$
Center Density $(k)$ $4.902\pm 0.413$ $\mathrm{stars\,arcmin^{-2}}$
Hill radius $28.608$ $\mathrm{arcmin}$
Gravitational bound radius $42.842$ $\mathrm{arcmin}$
@@
Proper Motion R.A. $2.542\pm 0.0094$ $\mathrm{mas\,yr^{-1}}$
Proper Motion DEC. $-1.713\pm 0.0086$ $\mathrm{mas\,yr^{-1}}$
Half-Mass Relaxation time $9.808$ $\mathrm{Myr}$
Radial Velocity $-0.67\pm 33.056$ $\mathrm{km\,s^{-1}}$
Concentration Parameter $(C)$ $3.042$
@@
$Y_{\text{frac}}$ $0.18$
Minimum Segregation Time $1.364$ $\mathrm{Myr}$
Table 2: Summary of derived parameters for NGC 6383, focusing on members with at least a 60 percent membership probability. Key measurements include distances assessed through parallax and distance modulus, mean age and stellar formation range, metallicity, and total stellar membership. The table also outlines structural attributes such as core and tidal radii, absorption, and galactocentric distance, complemented by dynamic properties like proper motions, radial velocities, concentration parameters, specific radius measurements, the number of identified Young Stellar Objects (YSOs), the fraction of YSOs, and the minimum segregation time.
Figure 7: CMD of NGC 6383, depicting the classification of cluster members and their evolutionary states, determined using the Sagitta. The mean fit isochrone is denoted by the dashed black curve, while the median fit is shown with a dotted black line. Red stars indicate PMS Members with membership probabilities and PMS probabilities $\geq 0.6$. Orange diamonds represent PMS Probable Members with probabilities and PMS probabilities $\leq 0.6$. Blue stars signify MS Members with membership probabilities $\leq 0.6$ and PMS probabilities $<0.6$, and blue-violet diamonds are for MS Probable Members with probabilities $>0.6$ and PMS probabilities $<0.6$. The shaded area around the isochrone visualizes the uncertainty within the parameter fits, represented by a normal distribution. Additionally, error bars, plotted at regular intervals along the color axis, mean color error.
@@ 2.2 AsTeCA
AsTeCA, or Automated Stellar Cluster Analysis (Perren et al., 2015), is an open-source Python tool designed to automate standard tests for determining basic parameters of stellar clusters. For accurate estimation of a cluster’s metallicity, age, and extinction values, we employed AsTeCA’s isochrone fitting process. We utilized the MIST isochrones (Dotter, 2016; Choi et al., 2016) along with the photometric systems of Gaia EDR3 and 2MASS. The Initial Mass Function (IMF) settings followed Chabrier et al. (2014), with fixed values for $\alpha=0.09$, $\beta=0.94$, and differential reddening set to zero.
The priors for metallicity $Z$, logarithmic age, and absorption $A_{V}$ were defined as uniform distributions ranging from $[0.01,0.045]$, $[5,10.3]$, and $[0.5,2]$, respectively, aligned with findings from Pulgar-Escobar & Henríquez-Salgado (2024). The distance modulus was modeled as $\mathcal{N}(10.3,0.2)$, with the mean value following the expected distance modulus obtained from the parallax. We opted for the median solution and chose to utilize the No-U-Turn Sampler (NUTS) from PyMC, prioritizing its efficiency over the Approximate Bayesian Computation (ABC) suggested by ASteCA.
Figure 8: Left panel: Color-Magnitude Diagram (CMD) for NGC 6383 showing $G_{\text{mag}}$ against $G_{\text{BP}}-G_{\text{RP}}$ with isochrones for $\log(\text{age})$ ranging from 6.2 to 7.0. Middle panel: CMD using $G_{\text{mag}}$ versus $G_{\text{RP}}-J$. Right panel: Color-Color diagram $G_{\text{RP}}-G_{\text{RP}}$ against $G_{\text{RP}}-J$ with the equivalent color-color isochrones. Red stars represent PMS Members, orange diamonds indicate PMS Probable Members, blue stars are MS Members, and blue-violet diamonds show MS Probable Members, as in Fig. 7. Isochrones are color-coded to represent different ages, highlighting the evolutionary progression of members within NGC 6383. Each plot includes a legend indicating the initial metallicity ($Z_{\text{ini}}=0.023$).

### v2

@@ 2.1.1 Membership determination
To identify potential members of NGC 6383, we employed the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (Campello et al., 2013; McInnes et al., 2017), known for its effectiveness in identifying clusters of varying shapes, densities, and sizes without requiring a predetermined number of clusters (Hunt & Reffert, 2021), which makes HDBSCAN an ideal choice for this analysis, offering significant benefits over traditional clustering algorithms, especially in terms of robustness against noisy data and outliers.
@@
After narrowing the data to a final subset of $15276$ sources, HDBSCAN was applied, focusing on proper motions as the primary clustering parameters.
@@
To refine our selection of hyperparameters for optimal cluster identification, we utilized a systematic approach. This involved the iterative testing of various minimum cluster sizes to find the setup that enhances cluster size while ensuring high cluster separation strength, as indicated by the maximum $\lambda$ value. Our strategy aimed at selecting leaf clusters to generate numerous small, homogeneous groups, with the goal of including the maximum number of stars without compromising the separation strength. While this approach risks increasing false positives, it is essential for capturing all potential cluster sources, particularly for studying tidal sources. To mitigate this risk, we derived a pseudo-probability metric. This metric was calculated by determining the number of times a source was identified as a cluster member across multiple iterations, divided by the total number of iterations, and then multiplying this value by the HDBSCAN membership probability Sources with a resulting probability greater than $0.5$ were retained. We consider sources with a membership probability between 0.6 and 0.8 as probable members, while those with a probability greater than 0.8 are classified as members. The variation in cluster sizes and the threshold for minimum cluster size are plotted in Fig. 16. Additionally, we used a Euclidean metric for clustering based on proper motions, finding that while the Haversine metric was tested, it produced non-physical clusters.
@@
Finally, we used the AstroPy Sigma Clipping (Astropy Collaboration et al., 2013, 2018, 2022) utility to apply a $2\sigma$ outlier rejection around the mode of the parallax distribution in order to obtain the final sample of likely members of the cluster. This step provided the dataset of cluster members.
Figure 5: Radial density profile of NGC 6383 shown in logarithmic scale. The plot overlays the observed stellar density (blue points) against the King model (solid fuchsia line). Dashed red and dotted red lines mark the core radius ($R_{c}=1.95\,\text{arcmin}$) and the tidal radius ($R_{t}=40.4\,\text{arcmin}$), respectively. The dash-dot green line represents the $b$ parameter of the background level. The shaded area in blue indicates the $1\sigma$ uncertainty range of the King profile.
Figure 6: Spatial distribution of probable members and members overlaid on a DSS2-red image, centered on the derived cluster center marked by a blue cross. The image details various structural and dynamical boundaries of the cluster as follows: the tidal radius $R_{t}$ is depicted with a dashed dark slate gray line, the core radius $R_{c}$ with a solid dark blue line, the half-light radius $R_{\mathrm{hl}}$ with a dash-dot burgundy line, the half-mass radius $R_{\mathrm{hm}}$ with a solid magenta line, the Hill radius $R_{\mathrm{hill}}$ with a dashed yellow line, and the gravitational bound radius $R_{\mathrm{bound}}$ with a solid orange line. A scale bar indicating $5\,\mathrm{pc}$ is positioned at the bottom right to provide spatial reference relative to the cluster scale.
@@ 2.1.2 Parallax and Distance Estimation
Distance estimation from parallax is challenging due to measurement errors, making the conversion from parallax to distance via the simple inverse method unreliable (Bailer-Jones, 2015; Bailer-Jones et al., 2021). Choosing an appropriate prior distribution is crucial. When analyzing the parallax measurements of stars within an open cluster, and assuming a simplified one-dimensional perspective, we assume that the parallax values follow a normal distribution. Additionally, we assume a gamma distribution for the distribution of distances.
To address this challenge, we obtained photo-geometric and geometric distances from Bailer-Jones et al. (2021), who introduced a Bayesian approach incorporating a prior distribution tailored to different regions of the sky, improving distance estimates from Gaia EDR3 data. Geometric distances are derived using only parallax and prior information, while photogeometric distances combine parallax with photometric data to refine the distance estimate. However, we opted not to use photogeometric distances due to the presence of PMS stars and gas in the cluster, which could distort the magnitudes and colors.
In our hierarchical model, we calculate the mean $(\mu_{\mathrm{prior}})$ between the frequentist average distance derived from parallaxes and the distances, using $\mathcal{U}(0.5\mu_{\mathrm{prior}},1.5\mu_{\mathrm{prior}})$ as the prior for the cluster’s mean distance. Our analysis is limited to members with fractional parallax errors below $0.1$ to ensure accuracy.
@@ 2.1.3 Proper motions
We used a two-dimensional Gaussian to model the distribution of the proper motions of the cluster members. A normal prior distribution was assigned to the mean proper motions $\left[\overline{\mu_{\alpha}},\overline{\mu_{\delta}}\right]$, based on frequentist means and standard deviations. For the standard deviations, we used half-Normal priors with a width equal to the frequentist standard deviation of the proper motions, and the correlation coefficient between the proper motions was uniformly distributed over the interval $[-1,1]$. Additionally, we estimated the mean projected velocity of the cluster, defined as the quadrature sum of the proper motion components.
Table 2: Summary of derived parameters for NGC 6383, focusing on sources with at least a 60 percent membership probability. Key measurements include distances assessed through parallax and distance modulus, mean age and stellar formation age, metallicity, and total stellar membership. The table also includes structural parameters such as core and tidal radii, absorption, and galactocentric distance, complemented by dynamic properties like proper motions, radial velocities, concentration parameters, specific radius measurements, the number of identified young stellar objects (YSOs), the fraction of YSOs, and the minimum segregation time.
@@
Distance $(\varpi)$ $1.11\pm 0.04$ $\mathrm{kpc}$
Distance (D.M.) $1.15_{-0.13}^{+0.15}$ $\mathrm{kpc}$
Distance modulus $10.3\pm 0.3$ $\mathrm{mag}$
Age $3.53^{+1.40}_{-1.00}$ $\mathrm{Myr}$
Stellar formation range $1.58-6.31$ $\mathrm{Myr}$
Metallicity ($Z$) $0.024\pm 0.008$
@@
Absorption $(A_{V})$ $1.24\pm 0.26$ $\mathrm{mag}$
Galactocentric distance $R_{\mathrm{GC}}$ $7.190\pm 0.004$ $\mathrm{kpc}$
Core radius $(R_{c})$ $1.95\pm 0.19$ $\mathrm{arcmin}$
Background $(b)$ $0.011\pm 0.006$ $\mathrm{stars\,arcmin^{-2}}$
Tidal radius $(R_{t})$ $40.4\pm 14.3$ $\mathrm{arcmin}$
Center density $(k)$ $4.91\pm 0.44$ $\mathrm{stars\,arcmin^{-2}}$
Hill radius $28.3\pm 1.0$ $\mathrm{arcmin}$
Gravitational bound radius $42.8\pm 1.6$ $\mathrm{arcmin}$
@@
Proper motion R.A. $2.540\pm 0.009$ $\mathrm{mas\,yr^{-1}}$
Proper motion Dec. $-1.710\pm 0.009$ $\mathrm{mas\,yr^{-1}}$
Half-mass relaxation time $13.6\pm 3.9$ $\mathrm{Myr}$
Radial velocity $-15.1\pm 31.1$ $\mathrm{km\,s^{-1}}$
Concentration parameter $(C)$ $3.03$
@@
$Y_{\text{frac}}$ $0.28^{+0.06}_{-0.06}$
Minimum segregation time $1.80\pm 0.42$ $\mathrm{Myr}$
Figure 7: CMD of NGC 6383 depicting the classification of cluster members and their evolutionary states, determined using Sagitta. The best-fit isochrones are represented by three lines: the dashed black curve for the mean fit, the dotted black line for the median fit, and the long dashed with offset line for the mode fit. Red stars indicate PMS Members with PMS probabilities $\geq 0.6$ and membership probabilities $\geq 0.8$. Orange diamonds represent PMS Probable Members with PMS probabilities $\geq 0.6$ and membership probabilities $<0.8$. Blue stars show non-PMS Members, which are sources with PMS probabilities $<0.6$ and membership probabilities $\geq 0.8$, indicating that they do not have enough probability to be classified as PMS, but they are not necessarily confirmed MS stars. Blue-violet diamonds represent non-PMS Probable Members, with PMS probabilities $<0.6$ and membership probabilities $<0.8$. Black stars and diamonds indicate Members and Probable Members with unavailable 2MASS data. The shaded area around the isochrone visualizes the uncertainty in the parameter fits, represented by a normal distribution. Additionally, error bars are plotted at regular intervals along the color axis to depict the median magnitude and color error.
Figure 8: Left panel: CMD of NGC 6383 showing $G_{\text{mag}}$ against $G_{\text{BP}}-G_{\text{RP}}$ with isochrones for $\log(\mathrm{age}\leavevmode\nobreak\ \mathrm{yr}^{-1})$ ranging from $6.20$ to $7.00$. Middle panel: CMD using $G_{\text{mag}}$ versus $G_{\text{RP}}-J$. Right panel: Color-Color diagram $G_{\text{RP}}-G_{\text{RP}}$ against $G_{\text{RP}}-J$ with the equivalent color-color isochrones. The symbol notation follows that of Fig. 7, representing different member classifications and evolutionary states within NGC 6383. Isochrones are color-coded to represent different ages, highlighting the evolutionary progression of members within NGC 6383. Each plot includes a legend indicating the initial metallicity ($Z_{\text{ini}}=0.024$).
2.1.4 Center determination
To determine the cluster’s center, we used a weighted Kernel Density Estimation (KDE) from the Python library sci-kit learn (Pedregosa et al., 2011), assigning weights inversely proportional to the distance from the mean proper motion. Optimal KDE parameters were determined through Grid Search Cross-Validation within the package, exploring a range of bandwidths—from the mean positional error to the search cone radius—and all kernel types. The location of maximum density was taken as the cluster’s center.
2.1.5 Radial density profile
To determine the structural parameters of the cluster, we utilized the King (1962) density profile. The numerical density $\rho$, was computed by dividing the cluster area into concentric annuli, each containing an equal number of stars. The number of annuli, $K$, adheres to the equiprobable bin rule ($K=2n^{2/5}$), where $n$ is the star count within the cluster. Fitting the King model to the density data allows for estimating the posterior distribution of the structural parameters.
Priors for the model parameters were set as follows: $b\sim\mathcal{U}(0,2\rho_{\mathrm{min}})$, $k\sim\mathcal{U}(0,2\rho_{\mathrm{max}})$, $R_{c}\sim\mathcal{U}(0,0.8R_{t})$, and $R_{t}\sim\mathcal{U}(R_{c},1.5T_{\mathrm{max}})$, where $T_{\mathrm{max}}$ represents the highest value between the Hill radius and the gravitational bound radius. This ensures that the chosen parameter space for $R_{c}$ and $R_{t}$ is within astrophysical valid limits.
The Hill Radius $(R_{\mathrm{hill}})$, which defines the region around a star cluster where its gravitational influence dominates over that of the Galaxy, can be calculated following Carrera et al. (2019):
where $R_{\mathrm{gc}}=7.19\pm 0.004\,\mathrm{kpc}$ denotes the galactocentric distance, $m_{c}$ the cluster mass from Hunt & Reffert (2024), with a value of $902.27\pm 92.3\leavevmode\nobreak\ \mathrm{M_{\odot}}$, and $M_{\mathrm{gc}}=1.43\times 10^{11}\pm 9.71\times 10^{7}\,\leavevmode\nobreak\ \mathrm{M_{\odot}}$ the Galactic mass within the cluster’s orbit, calculated as
$M_{\mathrm{gc}}=2\times 10^{8}\,M_{\odot}\left(\frac{R_{\mathrm{gc}}}{30\,\mathrm{pc}}\right)^{1.20}.$ (2)
$r_{\mathrm{bound}}=\left(\frac{Gm_{\mathrm{c}}}{2(A-B)^{2}}\right)^{1/3},$ (3)
@@ 2.2 ASteCA
Automated Stellar Cluster Analysis (ASteCA) (Perren et al., 2015), is an open-source Python tool designed to automate standard tests for determining basic parameters of stellar clusters. For accurate estimation of a cluster’s metallicity, age, and extinction values, we employed ASteCA’s isochrone fitting process. We utilized the MIST isochrones (Dotter, 2016; Choi et al., 2016) along with the photometric systems of Gaia EDR3 and 2MASS. The Initial Mass Function (IMF) settings followed Chabrier et al. (2014), with fixed values for $\alpha=0.090$, $\beta=0.940$, and differential reddening set to zero.
The priors for metallicity $Z$, logarithmic age, and absorption $A_{V}$ were defined as uniform distributions ranging from $[0.001,0.045]$, $[6.00,7.00]$, and $[0.5,2]$, respectively, aligned with findings from Pulgar-Escobar & Henríquez-Salgado (2024). The distance modulus was modeled as $\mathcal{N}(10.3,0.2)$, with the mean value corresponding to the expected distance modulus obtained from the parallax. We chose to use the mode of the posterior distribution for our analysis and utilized the NUTS from PyMC, prioritizing its efficiency over the Approximate Bayesian Computation (ABC) suggested by ASteCA.
Figure 9: Radial velocity plotted against G-band magnitude for NGC 6383 sources categorized by the presence of robust radial velocity amplitude data from Gaia. This measure reflects the total amplitude in the radial velocity time series following outlier removal. Yellow circles indicate sources without amplitude measurements, displaying radial velocity uncertainties with error bars. Blue circles represent sources with amplitude data, color-coded by amplitude value according to the scale on the right.
Figure 10: Histograms of the parameters obtained with Sagitta for NGC 6383 probable members, including all data, members with a PMS probability over $60\%$, members with a PMS probability under $60\%$, and stars with missing 2MASS data. The upper panel shows the distribution of PMS probability values, indicating the likelihood of stars being PMS. The middle panel presents the distribution of logarithmic age values, illustrating the range and frequency of stellar ages within the cluster. The lower panel displays the distribution of visual extinction values ($A_{V}$), representing the amount of dust extinction affecting the observed stars.